# Safe Pareto Improvements for Expected Utility Maximizers in Program Games

### v4

@@ 4.3 Guaranteeing PMM payoffs using CSR
${\widetilde{\mathbf{RN}}}^{i}(\mathbf{RN}^{-i},\boldsymbol{a}({\widetilde{\textbf{p}}^{i^{\textnormal{def}}}}))$ $=\mathbf{RN}^{i}(\mathbf{RN}^{-i},\boldsymbol{a}({\widetilde{\textbf{p}}^{i^{\textnormal{def}}}}))\cup\mathbf{PMP}_{i}(\mathbf{ro}(\textbf{p},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}}))).$
@@
    Suppose either $p_{i}$ is in a subjective equilibrium of $G(\mathcal{P}\cup\bigtimes_{i=1}^{{n}}P_{i}^{\textnormal{RN}})$, or, for some player $j$, $p_{i}$ is an element of some $\textbf{p}_{-j}$ in the support of $\beta_{j}$. Suppose $p_{i}\notin P_{i}^{\textnormal{RN}}$. Then for any $\textbf{p}_{-i}\in\bigtimes_{j\neq i}P_{j}^{\textnormal{RN}}$, we have $p_{i}({\textbf{p}}^{\textnormal{def}}_{-i})=p_{i}(\textbf{p}_{-i})$.
@@
    Let $\textbf{p}_{-i}\in\bigtimes_{j\neq i}P_{j}^{\textnormal{RN}}(\mathbf{RN}^{j})$ be in the support of $\beta_{i}$, and take any $p_{i}\in P_{i}^{\textnormal{RN}}(\mathbf{RN}^{i})$ with PMP-extension $\widetilde{p_{i}}$. For all $j\neq i$ and all $\boldsymbol{a}$, we have $\mathbf{RN}^{j}(\mathbf{RN}^{-j}_{{\widetilde{\mathbf{RN}}}^{i}},\boldsymbol{a})=\mathbf{RN}^{j}(\mathbf{RN}^{-j},\boldsymbol{a})\cup{V}$ for some ${V}\subseteq\mathbf{PMP}_{i}(\mathbf{ro}(\textbf{p},\boldsymbol{a}))$.
$x$$y$
$\boldsymbol{a}({\textbf{p}}^{\textnormal{def}})$
$\mathbf{ro}(\textbf{p},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}}))$
$\mathbf{PMP}_{1}(\mathbf{ro}(\textbf{p},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}})))$
Figure 4: Illustration of the argument for Theorem 3, for ${n}=2$ players. By default, the renegotiation outcome is the black circle, $\mathbf{ro}(\textbf{p},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}}))$. Player 1 considers whether to add to their renegotiation set $\mathbf{RN}^{1}(\mathbf{RN}^{2},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}}))$ their PMP of the black circle, the black striped segment $\mathbf{PMP}_{1}(\mathbf{ro}(\textbf{p},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}})))$. Player 1 is certain that player 2 would not change their set $\mathbf{RN}^{2}(\mathbf{RN}^{1},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}}))$ in response to this addition in a way that would make player 1 worse off (Assumption 9). This is because the only change player 1 has made is to add outcomes that make both players weakly better off than $\mathbf{ro}(\textbf{p},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}}))$ and do not make player 2 strictly better off.
Theorem 3.
Let $G(\mathcal{P})$ be a program game, and $\mathbf{sel}$ be any transitive selection function. Suppose the action sets of $G$ are continuous, so that for any $\boldsymbol{a}\in\mathcal{A}$, player $i$’s PMP of that action profile $\mathbf{PMP}_{i}(\boldsymbol{a})$ is nonempty. [^11] This is practically without loss of generality, because our program game setting can be extended to a setting where players can use correlated randomization (see, e.g., Kalai et al. (2010)). Let $\boldsymbol{\beta}$ be any belief profile satisfying the assumption that players are (i) certain that CSR won’t be punished and (ii) certain that PMP-extension won’t be punished (Assumption 9).
Then, for any subjective equilibrium $(\textbf{p},\boldsymbol{\beta})$ of $G(\mathcal{P}\cup\bigtimes_{i=1}^{{n}}P_{i}^{\textnormal{RN}})$ where $U_{i}(\textbf{p})<u_{i}^{\textnormal{PMM}}$ for some $i$, there exists $\textbf{p}^{\prime}$ such that:
  1. 1.
    For all $i$, $p_{i}^{\prime}$ is the PMP-extension of $p_{i}$.
  2. 2.
    $\textbf{U}(\textbf{p}^{\prime})\succeq\textbf{u}^{\textnormal{PMM}}$.
  3. 3.
    $(\textbf{p}^{\prime},\boldsymbol{\beta})$ is a subjective equilibrium of $G(\mathcal{P}\cup\bigtimes_{i=1}^{{n}}P_{i}^{\textnormal{RN}})$.
Proof Sketch.
Assumption 9i implies that players always use CSR programs. Consider any renegotiation outcome $\boldsymbol{a}$ worse for some player than the PMM, which is achieved by player $i$’s “old” program against some counterpart. By Assumption 9ii, no other player $j$ punishes $i$ for adding their PMP of that outcome, $\mathbf{PMP}_{i}(\boldsymbol{a})$, to their renegotiation set (in their “new” program). So the renegotiation outcome of the new program against $j$ is only different from that of the old program if $j$ is also willing to renegotiate to some outcome in $\mathbf{PMP}_{i}(\boldsymbol{a})$. But in that case, because the selection function is transitive, the new renegotiation outcome is no worse for $i$ than under the old program. Therefore, each player always prefers to replace a given program with its PMP-extension, and when all players use PMP-extended programs, the Pareto frontier of their agreement set only includes outcomes guaranteeing each player their PMM payoff. $\square$
Remark: Notice that the argument above does not require that players refrain from using programs that implement other kinds of SPIs, besides PMP-extensions.
First, the PMP-extension can be constructed from any default program, including, e.g., a CSR program whose renegotiation set is only extended to include the player’s most-preferred outcome, not their PMP (informally, call this a “self-favoring extension”). And even if a player uses a self-favoring extension as their final choice of program (the “outer loop”), they are still incentivized to use the PMP-extension within their default program (“inner loop”).
Second, while it is true that an analogous argument to the proof of Theorem 3 could show that a player is weakly better off (in expected utility) using a self-favoring extension than not extending their renegotiation set at all, this does not undermine our argument. This is because, as we claimed at the start of this section, it is reasonable to assume that among programs with equal expected utility, players prefer a program with a larger renegotiation set — i.e., prefer to include their PMP. On the other hand, this same argument doesn’t imply a player will include outcomes that Pareto-dominate the PMM as well, because those outcomes will be worse for that player and better for some counterpart than the player’s most-preferred outcome (i.e., it’s plausible that the non-punishment assumption would be violated).
We can now formalize the claim that CSR is an SPI that partially solves SPI selection: The mapping from programs p to instances of Algorithm 2 with these programs as defaults, for any profile $(\mathbf{RN}^{i})_{i=1}^{{n}}$ used in subjective equilibrium under the assumptions of Theorem 3, is an SPI. And, this SPI guarantees players their PMM payoffs.
Proposition 4.
For $i=1,\dots,{{n}}$, for some selection function $\mathbf{sel}$, define $f_{i}^{\mathbf{RN}^{i}}:P_{i}\rightarrow P_{i}^{\textnormal{RN}}(\mathbf{RN}^{i})$ such that, for each $p_{i}\in P_{i}$, $f_{i}^{\mathbf{RN}^{i}}(p_{i})$ is of the form given in Algorithm 2 with ${f_{i}^{\mathbf{RN}^{i}}(p_{i})}^{\textnormal{def}}=p_{i}$. Then, under the assumptions of Theorem 3, for any $(\mathbf{RN}^{i})_{i=1}^{{n}},{\textbf{p}}^{\textnormal{def}}$ such that for all $\tilde{\boldsymbol{a}},\mathbf{RN}^{-i}$, $\mathbf{PMP}_{i}(\mathbf{ro}(\textbf{p},\tilde{\boldsymbol{a}}))\subseteq\mathbf{RN}^{i}(\mathbf{RN}^{-i},\tilde{\boldsymbol{a}})$:
  1. 1.
    The function $\textbf{f}^{\mathbf{RN}}:\textbf{p}\mapsto(f_{i}^{\mathbf{RN}^{i}}(p_{i}))_{i=1}^{{n}}$ is an SPI.
  2. 2.
@@
This follows immediately from the argument used to prove Theorem 3. ∎
@@
4.4 Iterated CSR and tightness of the PMM bound
So far, we have considered renegotiation that takes place in one round. We might expect, however, that if players renegotiate for multiple rounds indefinitely, and they are required to take a strict Pareto improvement at every round of renegotiation, they are guaranteed payoffs that nontrivially exceed the PMM. As we will show, this is not always true.
Then, Proposition 5 shows that our PMM payoff bound is tight in bargaining problems, that is, games where no more than one player can achieve their best feasible payoff. The intuition for this result is given in Example 4.3.
@@ 4.4 Iterated CSR and tightness of the PMM bound
Example 4.3.
5 Discussion
Using renegotiation to construct SPIs in program games is a rich and novel area, with many directions to explore. To name a few: What are necessary and sufficient conditions on players’ beliefs for conditional set-valued renegotiation to guarantee efficiency? What do unilateral SPIs (Oesterheld and Conitzer, 2022) look like in this setting? When are SPIs used in sequential, rather than simultaneous-move, settings? And how can we use this theory to build real-world AI systems that avoid the harms of bargaining failure?
Acknowledgments
We thank Caspar Oesterheld, Lukas Finnveden, Vojtech Kovarik, and Alex Kastner for helpful comments on this paper.
Appendix A Notation Table
Table 2: Key notation
Symbol Description
$\mathbf{rn}^{i}$ renegotiation function for player $i$ (maps an action profile to a Pareto-improved
action profile)
$\mathbf{RN}^{i}$ set-valued renegotiation function for player $i$ (maps other player’s set-valued
renegotiation function and an action profile to a set of Pareto-improved action profiles)
$P_{i}^{\textnormal{rn}}(\mathbf{rn}^{i})$ set of renegotiation programs (Algorithm 1) that use the renegotiation function $\mathbf{rn}^{i}$
$P_{i}^{\textnormal{RN}}(\mathbf{RN}^{i})$ set of conditional set-valued renegotiation programs (Algorithm 2) that use the
set-valued renegotiation function $\mathbf{RN}^{i}$
$P_{i}^{\textnormal{IRN}}(\mathbf{RN}^{i}({K}))$ set of iterated conditional set-valued renegotiation programs (Algorithm 3) that use
the ${K}$-tuple of set-valued renegotiation functions $\mathbf{RN}^{i}({K})$
$p_{i}^{\textnormal{def}}$ default program for a program $p_{i}$ in $P_{i}^{\textnormal{rn}}$, $P_{i}^{\textnormal{RN}}$, or $P_{i}^{\textnormal{IRN}}$
$\mathbf{sel}$ selection function (maps a set of action profiles to an action profile that is efficient
within that set)
AS agreement set (as a function of a focal player’s set-valued renegotiation function,
the profile of their counterparts’ set-valued renegotiation functions, and
a miscoordinated action profile)
$\textbf{u}^{\textnormal{PMM}}$ Pareto meet minimum
$\mathbf{PMP}_{i}$ Pareto meet projection for player $i$ (maps an action profile to a particular set of
Pareto-improved action profiles)
Appendix B Proof of Proposition 2
See 2
Proof.
Let $(\textbf{p}^{*},\boldsymbol{\beta})$ be a subjective equilibrium of $G(\mathcal{P}\cup\bigtimes_{i=1}^{{n}}P_{i}^{\textnormal{rn}})$, where for some $i$, $p_{i}^{*}\notin P_{i}^{\textnormal{rn}}$; and let $\mathbf{rn}$ be an arbitrary renegotiation function. Define $f_{i}^{*}:P_{i}\rightarrow P_{i}^{\textnormal{rn}}(\mathbf{rn})$ such that, for each $p_{i}\in P_{i}$, $f_{i}^{*}(p_{i})$ is of the form given in Algorithm 1 with ${f_{i}^{*}(p_{i})}^{\textnormal{def}}=p_{i}$, for the given $\mathbf{rn}$.
Let $\textbf{p}_{-i}$ be in the support of $\beta_{i}$. We will show that against any such $\textbf{p}_{-i}$, the new program $f_{i}^{*}(p_{i}^{*})\in P_{i}^{\textnormal{rn}}(\mathbf{rn})$ is weakly better for player $i$ than $p_{i}^{*}$ (hence $f_{i}^{*}(p_{i}^{*})$ is better in expectation).
Now, let $\textbf{p}^{\prime}$ be the program profile such that:
  • $p_{i}^{\prime}=p_{i}^{*}$ if $p_{i}^{*}\in P_{i}^{\textnormal{rn}}(\mathbf{rn}^{i})$ for some $\mathbf{rn}^{i}$
  • $p_{i}^{\prime}=f_{i}^{*}(p_{i}^{*})$ otherwise.
Clearly, then, (1) each $p_{i}^{\prime}\in P_{i}^{\textnormal{rn}}$. And since the above argument holds for any $i$, we have $f_{i}^{*}(p_{i}^{*})\in\operatorname*{arg\,max}_{p_{i}\in P_{i}\cup P_{i}^{\textnormal{rn}}}\mathbb{E}_{\textbf{p}_{-i}\sim\beta_{i}}U_{i}(\textbf{p}),$ i.e., (2) $(\textbf{p}^{\prime},\boldsymbol{\beta})$ is a subjective equilibrium. ∎
Appendix C Proof of Claim that Coordination on the Selection Function Does Not Change Players’ Payoffs
Appendix D Proof of Theorem 3
See 3
Proof.
  • ${p_{i}^{\prime\prime}}^{\textnormal{def}}=p_{i}^{\prime}$, and
  • ${\widetilde{p_{i}}}^{\textnormal{def}}=p_{i}^{\textnormal{def}}$, and
  • For all $\textbf{p}_{-i}$, we have
    ${\widetilde{\mathbf{RN}}}^{i}(\mathbf{RN}^{-i},\boldsymbol{a}({\widetilde{\textbf{p}}^{i^{\textnormal{def}}}}))=\mathbf{RN}^{i}(\mathbf{RN}^{-i},\boldsymbol{a}({\widetilde{\textbf{p}}^{i^{\textnormal{def}}}}))\cup\mathbf{PMP}_{i}(\mathbf{ro}(\textbf{p},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}}))).$
  • PMP of default renegotiation outcome not in a counterpart’s set: By construction $\boldsymbol{a}({\widetilde{\textbf{p}}^{i^{\textnormal{def}}}})=\boldsymbol{a}({\textbf{p}}^{\textnormal{def}})$. If for some $j\neq i$, player $i$’s PMP of the default renegotiation outcome isn’t in $j$’s renegotiation set (in response to player $i$’s modified renegotiation set), then adding that projection to $i$’s set doesn’t make a difference. That is, let ${S}^{\mathbf{PMP}_{i}}=\mathbf{PMP}_{i}(\mathbf{ro}(\textbf{p},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}})))\cap\bigcap_{j\neq i}\mathbf{RN}^{j}(\mathbf{RN}^{-j}_{{\widetilde{\mathbf{RN}}}^{i}},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}}))$. If ${S}^{\mathbf{PMP}_{i}}=\emptyset$, then by Assumption 9ii, $\textbf{AS}({\widetilde{\mathbf{RN}}}^{i},\mathbf{RN}^{-i},\boldsymbol{a}({\widetilde{\textbf{p}}^{i^{\textnormal{def}}}}))=\textbf{AS}(\mathbf{RN}^{i},\mathbf{RN}^{-i},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}}))$, so $\textbf{U}({\widetilde{\textbf{p}}^{i}})=\textbf{U}(\textbf{p})$.
  • If ${S}^{\mathbf{PMP}_{i}}\neq\emptyset$:
    • By definition, ${S}^{\mathbf{PMP}_{i}}\subseteq\mathbf{PMP}_{i}(\mathbf{ro}(\textbf{p},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}})))$. So, by Assumption 9ii, [^12] Remark: Notice that could be relaxed to: “Let $\textbf{p}_{-i}\in\bigtimes_{j\neq i}P_{j}^{\textnormal{RN}}(\mathbf{RN}^{j})$ be in the support of $\beta_{i}$, and let $p_{i}\in P_{i}^{\textnormal{RN}}(\mathbf{RN}^{i})$ be in a subjective equilibrium of $G(\mathcal{P}\cup\bigtimes_{i=1}^{{n}}P_{i}^{\textnormal{RN}})$ with PMP-extension $\widetilde{p_{i}}$. For all $j\neq i$ and all $\boldsymbol{a}$, we have $\mathbf{RN}^{j}(\mathbf{RN}^{-j}_{{\widetilde{\mathbf{RN}}}^{i}},\boldsymbol{a})=\mathbf{RN}^{j}(\mathbf{RN}^{-j},\boldsymbol{a})\cup{V}$ for some ${V}\subseteq\mathbf{PMP}_{i}(\mathbf{ro}(\textbf{p},\boldsymbol{a}))$.” We state the technically stronger assumption in the main text for simplicity. the only change to the agreement set due to the PMP-extension is that outcomes from player $i$’s PMP of the default renegotiation outcome are added: $\textbf{AS}({\widetilde{\mathbf{RN}}}^{i},\mathbf{RN}^{-i},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}}))=\textbf{AS}(\mathbf{RN}^{i},\mathbf{RN}^{-i},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}}))\cup{S}^{\mathbf{PMP}_{i}}$, where this set is nonempty. (Thus the new renegotiation outcome is $\mathbf{ro}({\widetilde{\textbf{p}}^{i}},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}}))=\mathbf{sel}(\textbf{AS}({\widetilde{\mathbf{RN}}}^{i},\mathbf{RN}^{-i},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}})))$.)
    • Lastly, we consider the new renegotiation outcome given the two cases for the default renegotiation outcome $\mathbf{ro}(\textbf{p},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}}))$:
      1. 1.
        Agreement achieved without PMP: Suppose $i$’s original program reached agreement with the other players, that is, we have $\mathbf{ro}(\textbf{p},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}}))=\mathbf{sel}(\textbf{AS}(\mathbf{RN}^{i},\mathbf{RN}^{-i},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}})))$. For all $\textbf{x}\in\mathbf{PMP}_{i}(\mathbf{ro}(\textbf{p},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}})))$, the following holds by the definition of PMP: $\textbf{u}(\textbf{x})\succeq\textbf{u}(\mathbf{ro}(\textbf{p},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}})))$. Since $\mathbf{sel}$ is a transitive selection function, therefore,
        $\textbf{u}(\mathbf{ro}({\widetilde{\textbf{p}}^{i}},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}})))\succeq\textbf{u}(\mathbf{ro}(\textbf{p},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}}))).$
      2. 2.
        No agreement without PMP: Otherwise, since by construction we have $\mathbf{ro}({\widetilde{\textbf{p}}^{i}},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}}))\in\mathbf{PMP}_{i}(\mathbf{ro}(\textbf{p},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}})))$, then
        $\textbf{u}(\mathbf{ro}({\widetilde{\textbf{p}}^{i}},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}})))\succeq\textbf{u}(\mathbf{ro}(\textbf{p},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}}))).$
Thus, for all $\textbf{p}_{-i}$ we have $U_{i}({\widetilde{\textbf{p}}^{i}})\geq U_{i}(\textbf{p})$. Applying the same argument for each player $j$, and with $\widetilde{\textbf{p}}=(\widetilde{p_{i}})_{i=1}^{{n}}$, it follows that $(\widetilde{\textbf{p}},\boldsymbol{\beta})$ is a subjective equilibrium. Now, assume this profile does not guarantee the PMM, that is, for some $i^{\prime}$, $u_{i^{\prime}}(\mathbf{ro}(\widetilde{\textbf{p}},\boldsymbol{a}({\widetilde{\textbf{p}}}^{\textnormal{def}})))<u_{i^{\prime}}^{\textnormal{PMM}}$. Let ${\widetilde{\mathbf{RN}}}^{-i}=({\widetilde{\mathbf{RN}}}^{j})_{j\neq i}$. Since $\mathbf{PMP}_{i}(\mathbf{ro}(\widetilde{\textbf{p}},\boldsymbol{a}({\widetilde{\textbf{p}}}^{\textnormal{def}})))\subseteq{\widetilde{\mathbf{RN}}}^{i}({\widetilde{\mathbf{RN}}}^{-i},\boldsymbol{a}({\widetilde{\textbf{p}}}^{\textnormal{def}}))$ for all $i$, it follows from the argument above that the players’ renegotiation sets have nonempty intersection, and that $\mathbf{sel}(\textbf{AS}({\widetilde{\mathbf{RN}}}^{i},{\widetilde{\mathbf{RN}}}^{-i},\boldsymbol{a}({\widetilde{\textbf{p}}}^{\textnormal{def}})))\in\bigcap_{i=1}^{{n}}\mathbf{PMP}_{i}(\mathbf{ro}(\widetilde{\textbf{p}},\boldsymbol{a}({\widetilde{\textbf{p}}}^{\textnormal{def}})))$ guarantees each player at least $u_{i}^{\textnormal{PMM}}$. This contradicts the assumption that $\widetilde{\textbf{p}}$ doesn’t guarantee the PMM, so $\textbf{U}(\widetilde{\textbf{p}})\succeq\textbf{u}^{\textnormal{PMM}}$ as required. ∎
Appendix E Proof of Proposition 5
See 5
@@ Appendix E Proof of Proposition 5
Checking that this subjective equilibrium satisfies the assumption of Theorem 3: Each player $i$’s beliefs put probability 1 on the other players using programs whose renegotiation sets $\mathbf{RN}^{j,(k)}$ are independent of $\mathbf{RN}^{i,(k)}$. Thus such a program does not respond differently to $p_{i}$ and $\widetilde{p_{i}}$ as defined in Definition 8. ∎

### v5

@@ Appendix C Proof of Theorem 9
${\widetilde{\mathbf{RN}}}^{i}(\mathbf{RN}^{-i},\boldsymbol{a}({\widetilde{\textbf{p}}^{i^{\textnormal{def}}}}))$ $=\mathbf{RN}^{i}(\mathbf{RN}^{-i},\boldsymbol{a}({\widetilde{\textbf{p}}^{i^{\textnormal{def}}}}))\cup\mathbf{PMP}_{i}(\boldsymbol{a}(\textbf{p})).$
@@
    Suppose either $p_{i}$ is in a subjective equilibrium of $G(\mathcal{P}\cup\mathcal{P}^{\text{RN}})$, or, for some player $j$, $p_{i}$ is an element of some $\textbf{p}_{-j}$ in the support of $\beta_{j}$. Suppose $p_{i}\notin P_{i}^{\textnormal{RN}}$. Then for any $\textbf{p}_{-i}\in\bigtimes_{j\neq i}P_{j}^{\textnormal{RN}}$, we have $p_{i}({\textbf{p}}^{\textnormal{def}}_{-i})=p_{i}(\textbf{p}_{-i})$.
@@
    Let $\textbf{p}_{-i}\in\bigtimes_{j\neq i}P_{j}^{\textnormal{RN}}(\mathbf{RN}^{j})$ be in the support of $\beta_{i}$, and take any $p_{i}\in P_{i}^{\textnormal{RN}}(\mathbf{RN}^{i})$ with PMP-extension $\widetilde{p_{i}}$. For all $j\neq i$ and all $\boldsymbol{a}$, we have $\mathbf{RN}^{j}(\mathbf{RN}^{-j}_{{\widetilde{\mathbf{RN}}}^{i}},\boldsymbol{a})=\mathbf{RN}^{j}(\mathbf{RN}^{-j},\boldsymbol{a})\cup{V}$ for some ${V}\subseteq\mathbf{PMP}_{i}(\boldsymbol{a}(\textbf{p}))$.
See 9
@@
  • ${p_{i}^{\prime\prime}}^{\textnormal{def}}=p_{i}^{\prime}$, and
  • ${\widetilde{p_{i}}}^{\textnormal{def}}=p_{i}^{\textnormal{def}}$, and
  • For all $\textbf{p}_{-i}$, we have
    ${\widetilde{\mathbf{RN}}}^{i}(\mathbf{RN}^{-i},\boldsymbol{a}({\widetilde{\textbf{p}}^{i^{\textnormal{def}}}}))=\mathbf{RN}^{i}(\mathbf{RN}^{-i},\boldsymbol{a}({\widetilde{\textbf{p}}^{i^{\textnormal{def}}}}))\cup\mathbf{PMP}_{i}(\boldsymbol{a}(\textbf{p})).$
  • PMP of default renegotiation outcome not in a counterpart’s set: By construction $\boldsymbol{a}({\widetilde{\textbf{p}}^{i^{\textnormal{def}}}})=\boldsymbol{a}({\textbf{p}}^{\textnormal{def}})$. If for some $j\neq i$, player $i$’s PMP of the default renegotiation outcome isn’t in $j$’s renegotiation set (in response to player $i$’s modified renegotiation set), then adding that projection to $i$’s set doesn’t make a difference. That is, let ${S}^{\mathbf{PMP}_{i}}=\mathbf{PMP}_{i}(\boldsymbol{a}(\textbf{p}))\cap\bigcap_{j\neq i}\mathbf{RN}^{j}(\mathbf{RN}^{-j}_{{\widetilde{\mathbf{RN}}}^{i}},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}}))$. If ${S}^{\mathbf{PMP}_{i}}=\emptyset$, then by Assumption 7ii, $\textbf{AS}({\widetilde{\mathbf{RN}}}^{i},\mathbf{RN}^{-i},\boldsymbol{a}({\widetilde{\textbf{p}}^{i^{\textnormal{def}}}}))=\textbf{AS}(\mathbf{RN}^{i},\mathbf{RN}^{-i},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}}))$, so $\textbf{U}({\widetilde{\textbf{p}}^{i}})=\textbf{U}(\textbf{p})$.
  • If ${S}^{\mathbf{PMP}_{i}}\neq\emptyset$:
    • By definition, ${S}^{\mathbf{PMP}_{i}}\subseteq\mathbf{PMP}_{i}(\boldsymbol{a}(\textbf{p}))$. So, by Assumption 7ii, [^9] Remark: Notice that this could be relaxed to: “Let $\textbf{p}_{-i}\in\bigtimes_{j\neq i}P_{j}^{\textnormal{RN}}(\mathbf{RN}^{j})$ be in the support of $\beta_{i}$, and let $p_{i}\in P_{i}^{\textnormal{RN}}(\mathbf{RN}^{i})$ be in a subjective equilibrium of $G(\mathcal{P}\cup\bigtimes_{i=1}^{{n}}P_{i}^{\textnormal{RN}})$ with PMP-extension $\widetilde{p_{i}}$. For all $j\neq i$ and all $\boldsymbol{a}$, we have $\mathbf{RN}^{j}(\mathbf{RN}^{-j}_{{\widetilde{\mathbf{RN}}}^{i}},\boldsymbol{a})=\mathbf{RN}^{j}(\mathbf{RN}^{-j},\boldsymbol{a})\cup{V}$ for some ${V}\subseteq\mathbf{PMP}_{i}(\boldsymbol{a}(\textbf{p}))$.” We state the technically stronger assumption in the main text for simplicity. the only change to the agreement set due to the PMP-extension is that outcomes from player $i$’s PMP of the default renegotiation outcome are added: $\textbf{AS}({\widetilde{\mathbf{RN}}}^{i},\mathbf{RN}^{-i},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}}))=\textbf{AS}(\mathbf{RN}^{i},\mathbf{RN}^{-i},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}}))\cup{S}^{\mathbf{PMP}_{i}}$, where this set is nonempty. (Thus the new renegotiation outcome is $\boldsymbol{a}({\widetilde{\textbf{p}}^{i}})=\mathbf{sel}(\textbf{AS}({\widetilde{\mathbf{RN}}}^{i},\mathbf{RN}^{-i},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}})))$.)
    • Lastly, we consider the new renegotiation outcome given the two cases for the default renegotiation outcome $\boldsymbol{a}(\textbf{p})$:
      1. (1)
        Agreement achieved without PMP: Suppose $i$’s original program reached agreement with the other players, that is, we have $\boldsymbol{a}(\textbf{p})=\mathbf{sel}(\textbf{AS}(\mathbf{RN}^{i},\mathbf{RN}^{-i},\boldsymbol{a}({\textbf{p}}^{\textnormal{def}})))$. For all $\textbf{x}\in\mathbf{PMP}_{i}(\boldsymbol{a}(\textbf{p}))$, the following holds by the definition of PMP: $\textbf{u}(\textbf{x})\succeq\textbf{u}(\boldsymbol{a}(\textbf{p}))$. Since $\mathbf{sel}$ is a transitive selection function, therefore,
        $\textbf{u}(\boldsymbol{a}({\widetilde{\textbf{p}}^{i}}))\succeq\textbf{u}(\boldsymbol{a}(\textbf{p})).$
      2. (2)
        No agreement without PMP: Otherwise, since by construction we have $\boldsymbol{a}({\widetilde{\textbf{p}}^{i}})\in\mathbf{PMP}_{i}(\boldsymbol{a}(\textbf{p}))$, then
        $\textbf{u}(\boldsymbol{a}({\widetilde{\textbf{p}}^{i}}))\succeq\textbf{u}(\boldsymbol{a}(\textbf{p})).$
Thus, for all $\textbf{p}_{-i}$ we have $U_{i}({\widetilde{\textbf{p}}^{i}})\geq U_{i}(\textbf{p})$. Applying the same argument for each player $j$, and with $\widetilde{\textbf{p}}=(\widetilde{p_{i}})_{i=1}^{{n}}$, it follows that $(\widetilde{\textbf{p}},\boldsymbol{\beta})$ is a subjective equilibrium. Now, assume this profile does not guarantee the PMM, that is, for some $i^{\prime}$, $u_{i^{\prime}}(\boldsymbol{a}(\widetilde{\textbf{p}}))<u_{i^{\prime}}^{\textnormal{PMM}}$. Let ${\widetilde{\mathbf{RN}}}^{-i}=({\widetilde{\mathbf{RN}}}^{j})_{j\neq i}$. Since $\mathbf{PMP}_{i}(\boldsymbol{a}(\widetilde{\textbf{p}}))\subseteq{\widetilde{\mathbf{RN}}}^{i}({\widetilde{\mathbf{RN}}}^{-i},\boldsymbol{a}({\widetilde{\textbf{p}}}^{\textnormal{def}}))$ for all $i$, it follows from the argument above that the players’ renegotiation sets have nonempty intersection, and that $\mathbf{sel}(\textbf{AS}({\widetilde{\mathbf{RN}}}^{i},{\widetilde{\mathbf{RN}}}^{-i},\boldsymbol{a}({\widetilde{\textbf{p}}}^{\textnormal{def}})))\in\bigcap_{i=1}^{{n}}\mathbf{PMP}_{i}(\boldsymbol{a}(\widetilde{\textbf{p}}))$ guarantees each player at least $u_{i}^{\textnormal{PMM}}$. This contradicts the assumption that $\widetilde{\textbf{p}}$ doesn’t guarantee the PMM, so $\textbf{U}(\widetilde{\textbf{p}})\succeq\textbf{u}^{\textnormal{PMM}}$ as required. ∎
Appendix D Iterated CSR and Tightness of the PMM Bound
In the main text, we considered renegotiation that takes place in one round. We might expect, however, that if players renegotiate for multiple rounds indefinitely, and they are required to take a strict Pareto improvement at every round of renegotiation, they are guaranteed payoffs that nontrivially exceed the PMM. As we will show, this is not always true.
Then, Proposition 5 shows that our PMM payoff bound is tight in bargaining problems, that is, games where no more than one player can achieve their best feasible payoff. The intuition for this result is given in Example D.1.
@@ Appendix D Iterated CSR and Tightness of the PMM Bound
Example D.1.
@@
Checking that this subjective equilibrium satisfies the assumption of Theorem 9: Each player $i$’s beliefs put probability 1 on the other players using programs whose renegotiation sets $\mathbf{RN}^{j,(k)}$ are independent of $\mathbf{RN}^{i,(k)}$. Thus such a program does not respond differently to $p_{i}$ and $\widetilde{p_{i}}$ as defined in Definition 8.